# Part I: Nonlinear Codes

### v1

@@ VI Conclusion
We have defined the concept of de Bruijn array codes which generalizes the concept of a de Bruijn array. The main interesting question is whether the necessary conditions of Lemma 1 are also sufficient. In other words, given four integers, $k$, $t$, $n$, $m$, such that $\Delta=2^{nm-k-t}$, does there exists a $(2^{k},2^{t};n,m)$-DBAC with $\Delta$ codewords? We conjecture that such a code exists, i.e., the necessary conditions are sufficient.
If $\Delta=1$ the problem was completely solved by Paterson [42]. The direct constructions together with the merge-or-split method to merge codewords and the recursive construction solve some families of parameters. First, these constructions were designed for the case where a PF$(n,k)$ exists, i.e., when $k\leq n<2^{k}$, i.e., $n<2^{k}\leq 2^{n}$, but by the transposing the codeword and applying the recursive constructions other sets of parameters can be obtained.
To prove that the necessary conditions are indeed sufficient more constructions are required. These are left for future research. Indeed, the span of parameters for different families are so wide that it will probably take some time before this will ultimately by solved. Similarly to de Bruijn array codes, we can define pseudo-random array codes, a generalization for pseudo-random arrays. While for de Bruijn array codes most of the constructions have a combinatorial flavor, for pseudo-random array codes all the constructions will have an algebraic flavor. These arrays were considered in [11] and a comprehensive work on them will be presented in the follow-up paper [5].
@@
The description of joining codewords from the codes obtained from the construction of Theorem 8 leads to two open questions associated with ordering of codewords of a given code $\mathbb{C}$.
@@ Acknowledgement
Part of this work have been presented at the IEEE International Symposium on Information Theory, Athens, Greece,, July 2024. This research was supported in part by the Israeli Science Foundation grant no. 222/19.

### v2

@@ VII Conclusion and Future Work
We have defined the concept of a de Bruijn array code, which generalizes the concept of a de Bruijn array. Necessary conditions for the existence of such code were given in Lemma 1. The first goal in investigating these codes is to find whether the necessary conditions are also sufficient. In other words, given four integers, $k$, $t$, $n$, $m$, such that $\Delta=2^{nm-k-t}$, does there exist a $(2^{k},2^{t};n,m)$-DBAC with $\Delta$ codewords? We conjecture that such a code exists, i.e., the necessary conditions are sufficient.
Several direct and recursive constructions were presented. These constructions yield de Bruijn array codes with various parameters and several infinite families of such codes. It should take more careful analysis on the constructed codes of this work to characterize exactly which classes of codes are obtained and what is required to obtain in order to solve the conjecture. To prove that the necessary conditions are indeed sufficient more constructions are required. These are left for future research. Indeed, the range of possible parameters is so wide that it will probably take some time before the conjecture is resolved.
Analogously to de Bruijn array codes, we can define pseudo-random array codes, a generalization of pseudo-random arrays. MacWilliams and Sloane [33] defined pseudo-random arrays as follows.
Definition 4
A pseudo-random array ${\cal A}$ is an $r\times s$ doubly-periodic array such that each $n\times m$ nonzero matrix is contained exactly once as a window in the array. Moreover, if ${\cal A}$ is such an array, then ${\cal A}+{\cal A}^{\prime}$, where ${\cal A}^{\prime}$ is a nontrivial shift of ${\cal A}$ and the addition is performed bit-by-bit, is another nontrivial shift of ${\cal A}$. This is a shift-and-add-property.
The definition of pseudo-random arrays is generalized to pseudo-random array codes and will be considered in the follow-up paper [5]. They were also considered in a preliminary version that discussed these array codes [11]. In the current paper the constructions of DBACs are based on combinatorial techniques and the constructed codes are nonlinear. In the follow-up paper [5] the constructions of pseudo-random array codes will be based on algebraic techniques and the constructed codes will be linear.
@@
There are many directions to continue this research including questions on different types of codes. For example, joining codewords from the codes obtained from Construction 1 leads to two open questions associated with an ordering of codewords of a given code $\mathbb{C}$.
@@ Acknowledgement
Part of this work was presented at the IEEE International Symposium on Information Theory, Athens, Greece, July 2024. This research was supported in part by the Israeli Science Foundation grant no. 222/19. The author is indebted to two anonymous reviewers whose comments considerably improved the presentation of the paper.